OpenStudy (anonymous):

usind integration by substitution what is the integral of (2x+1)(x^2+x+1)^7

OpenStudy (anonymous):

intrinsic philosophy

OpenStudy (anonymous):

Again, make the higher order term = u and the lower as du, i.e., u = x^2 + x + 1 du = (2x + 1)dx

OpenStudy (anonymous):

then it will become a power rule integration of the form:\[\int\limits_{a}^{b}u^{7}du\]

OpenStudy (anonymous):

let u=x^2+x+1, du = 2x+1, so integral of 1/(u)^7du, so its' 1/8(u)^8, sub u back in u get 1/8(x^2+x+1)+c

OpenStudy (anonymous):

bmp, u mean 1/u^7 ?

OpenStudy (anonymous):

Nope, that's a multiplication. If it was a division, we would've gotten a natural log integration pattern.

OpenStudy (anonymous):

aa, i completely saw a division sign -.- i must be tired lol

OpenStudy (lgbasallote):

lol we're all seeing division signs @_@

OpenStudy (anonymous):

got some error with my division one 2... =/

OpenStudy (anonymous):

first time doing a ques. like this..can some 1 show me all the steps...this is where i am so far

OpenStudy (anonymous):

Your du term is incorrect, you most likely committed a minor mistake there. It should be du = (2x+1)dx, not otherwise. And, yes, the integral will be of the form\[\frac{u^{8}}{8} + c\]where u = x^2 + x + 1

OpenStudy (anonymous):

\[\int\limits_{}^{}(2x+1)(x ^{2}+x+1)^{7}dx\] u=\[x ^{2}+x+1\] du=2x+1 \[\int\limits_{}^{}u ^{7}du\] (1\[(1\div8)u ^{8}+c\] \[(1\div8)(x ^{2}+x+1)^{8}+c\]

OpenStudy (anonymous):